The sum of two fractions is always a fraction.
 The sum of two fractions is always a fraction.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

The given statement isTrue.
For example: consider two fractions 10/5 and 15/5.
Sum of two fractions = (10 + 15)/5

= 25/5
=5
= 5/1

A fraction in which there is no common factor, except 1, in its numerator and denominator is called a fraction in the simplest or lowest form.
When 2 fractions are added, the result in most cases will be a fraction p/q form, but in some case if it does happen to be just a integer, it can always be written with denominator 1 (hence p/q form)




Thanks